OpenStudy (anonymous):

At a conference, the organizer determined that 32% of the attendees were between the ages of 18 and 29. Since there were 216 attendees in this age range, there were total attendees at the conference. Submit

OpenStudy (anonymous):

Total at conference = 216/x = 32/100 Set the problem up as a ratio

OpenStudy (anonymous):

Mrs.Debbie are you a teacher?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you get the answer

OpenStudy (anonymous):

You could also divide 216/.32

OpenStudy (anonymous):

yes mam

OpenStudy (anonymous):

OK..great

OpenStudy (anonymous):

my mom cant help me do this and all my teachers are on break

OpenStudy (anonymous):

Good old winter break...I'm signed in off and on daily.

OpenStudy (anonymous):

do u need more help

OpenStudy (anonymous):

no mam

OpenStudy (anonymous):

k